6. Find the amplitude of each function below- if no amplitude exists, write NA. a) y sinx b) y = - 3 cos (x) c) y = 2 tan (x) answer all please
Answer #1


And y = 72 sinne amplitulle = 2 / (1-1) so amplitude = 7% (b) y=-3 (053 coste (111 I ampletude = 3 (c) y= = 2 fanleg 4anx El-y = tanke) y = Sant-se) = odd funtion - Wah x y 10) - Secky y = Ycos (2) yl-x) = cos(-x) = cose Odd funtion so only (b) y=tan(6) y= = cota y = cosa sina where sinu=o fin undefined anx: 2T - ot, so X= is the point C) y= seca y = cosa where cosa=0 fln9 - 240$ 2 34 น RA 3 SME 2. -SL น x-inter septs - * , * , * ) y inter seepts no asymptotes 2(b). y indesseets = none ze intersect 을 출 존 asymptodes + y이 y즈, -22, =지 =2T 32 = 여 기 MS

Please upvote this answer

If you have any doubts then comment below